OpenStudy (anonymous):

In baseball, the distance of the paths between each pair of consecutive bases is 90 ft and the paths form right angles. How far does the ball need to travel if it is thrown from home plate directly to second base? I know we need to find the height of the triangle using the Pythagorean Theorem, but my answer was wrong :( Help please?

OpenStudy (anonymous):

127.279221 feet

OpenStudy (phi):

you are asking for the length of a diagonal of a square with sides = 90 you can use pythagoras or you can use the idea that you have a 45-45-90 triangle.

OpenStudy (saifoo.khan):

This might help? http://www.dsr.wa.gov.au/assets/images/Diagrams/Layout-of-a-baseball-diamond.gif

OpenStudy (phi):

can you post how you did this problem ?

OpenStudy (anonymous):

I did h^2 + 45^2 = 90^2

OpenStudy (phi):



OpenStudy (phi):

I don't see where you got h^2 + 45^2 = 90^2 but look at the picture...

OpenStudy (anonymous):

I was thinking of a triangle instead of a square. So would it be 90^2 + 90^2 = h^2 ?

OpenStudy (phi):

yes and \[ h= \sqrt{90^2 + 90^2} = \sqrt{2\cdot 90^2}= 90 \sqrt{2}\]

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (phi):

Think over how you got the wrong answer, and what you would do differently to get the right answer.